1.If current assets are $33,000 and total assets are $224,000, what percentage of total assets are current assets? (Answer as a percentage to 1 decimal point with no % sign needed, e.g. xx.x or 12.3)

2.A company's net income before income tax is $72,000. Income tax for the year was $5,000. Dividend requirements for preferred stock are $20,000. There are 20,000 shares of common stock outstanding. Earnings per common stock are: (Answer to the nearest $0.01, w/ no $ sign needed, e.g. x.xx or 1.23)

3.Jack and Jill Corporation's year-end 20xx balance sheet lists current assets of $225,000, fixed assets of $965,000, current liabilities of $149,000, and long-term debt of $215,000. What is Jack and Jill's total stockholders' equity? (Answer in whole $s, with no $ sign or commas, e.g. xxxxxx, or 123000)

4.Billy Bob, Inc.'s 201x income statement lists the following income and expenses: EBIT = $838,000, interest expense = $90,000, and net income = $612,000. What are the 201x taxes reported on the income statement? (Answer in whole $s with no $ sign or commas needed, e.g. xxx000, or 123000)
Answer #1

1.Current assets as a percentage of total assets = $ 33,000 / $ 224,000 * 100 = 14.7

2. Earnings per common stock : 2.35.

Earnings available for common stockholders = $ ( 72,000 - 5,000 - 20,000 ) = $ 47,000

Earnings per common stock = $ 47,000 / 20,000 shares = $ 2.35

3. Jack and Jill's Stockholders Equity = 826000.

Stockholders' Equity = Total Assets - Total Liabilities = $ ( 225,000 + 965,000 ) - $ ( 149,000 + 215,000) = $ 826,000.

4. Taxes reported on the income statement : 136000.